# Language Models Do Not Have Human-Like Working Memory

### v2

@@ Abstract
While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they fundamentally lack a core aspect of human cognition: working memory. Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization. Without working memory, individuals may produce unreal conversations, exhibit self-contradiction, and struggle with tasks requiring mental reasoning. Existing evaluations using N-back or context-dependent tasks fail as they allow LLMs to exploit accessible context rather than retain latent information. We introduce three novel tasks—(1) Number Guessing, (2) Yes-No Deduction, and Math Magic—that isolate internal representation from external context. Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, highlighting LLMs’ inability to retain and manipulate latent information. Our work establishes a new benchmark for evaluating working memory in LLMs and identifies this deficit as a critical obstacle to artificial general intelligence. Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated strong performance across diverse domains, including legal advice Guha et al. (2023), education Wen et al. (2024), translation Jiao et al. (2023), and healthcare Yang et al. (2024b). Beyond task performance, recent studies have examined their intrinsic behaviors, often drawing parallels with human cognition, such as personality Huang et al. (2024b), emotion Huang et al. (2024a), empathy Sorin et al. (2024), theory-of-mind Liu et al. (2024b), and values Wang et al. (2024b).
@@
One such behavior is memory, which has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024). [^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ By mid-2025, xAI, Anthropic, and Google integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively. Sam Altman even emphasized that “people want memory,” identifying it as a central feature of the forthcoming GPT-6.555https://futurism.com/disastrous-gpt-5-sam-altman-hyping-up-gpt-6 On the research side, memory modules are also actively studied Zhang et al. (2025); Hatalis et al. (2023); Xu et al. (2025).
@@
What is memory? Atkinson & Shiffrin (1968) categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning Baddeley & Hitch (1974). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023). As a central aspect of human cognition, memory has been widely benchmarked Du et al. (2025). However, working memory remains relatively underexplored.
@@
Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al. (2024); Zhang et al. (2024). Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context. Unlike humans, who cannot revisit prior steps, LLMs can simply attend to earlier tokens within their context window. Fig. 2 illustrates this discrepancy. To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context. Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response. [^6] A possibility remains that the response was given by chance.
@@
We ask: Do LLMs possess human-like working memory beyond evaluations that merely test their ability to attend to visible information? AI models are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al. (2022); Yao et al. (2023), and (2) the latent space over activations Hao et al. (2024); Geiping et al. (2025). Our focus is on the second mode, where evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization. Success in this capacity could enhance reasoning without reliance on Chain-of-Thought (CoT), as it directly tests the model’s ability to maintain objects and concepts internally. Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
@@
The central challenge in designing such evaluations is epistemic: how can we demonstrate the absence of internal memory when we cannot directly observe a model’s mind? To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context. Our experiments span 17 state-of-the-art LLMs, both proprietary and open-source, and reveal a consistent inability to perform such tasks. The failure is uniform across model families (GPT Hurst et al. (2024), Qwen Yang et al. (2024a), DeepSeek Liu et al. (2024a); Guo et al. (2025), LLaMA Grattafiori et al. (2024)) and reasoning approaches (CoT Wei et al. (2022) or o1-like reasoning Jaech et al. (2024)).
2 Background
@@ 2.1 How Working Memory is Evaluated
Human working memory is typically assessed using behavioral paradigms that require individuals to maintain and update information over short intervals. Common examples include the digit span task Miller (1956), where participants recall sequences of numbers of increasing length, and the N-back task Kirchner (1958), which requires identifying whether the current stimulus matches one presented $N$ steps earlier. These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function.
In the context of LLMs, working memory has been widely discussed. However, most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al. (2023); Guo et al. (2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLMs using N-back tasks. As noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window rather than actively maintaining information.
@@ 2.2 Existing Engineering Solutions
A separate line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al. (2024); Zeng et al. (2024). For example, Wang et al. (2024a) incorporate symbolic working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. Other approaches implement scratch spaces Lanchantin et al. (2023) or internal CoT mechanisms Jaech et al. (2024), which may partially mitigate working memory limitations identified in this paper. However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects. The distinction can be illustrated with an analogy: modern calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division. The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities. Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model itself has the basic cognitive capability of working memory. Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
@@ 3 Number Guessing
(a) Probabilities for OpenAI GPT models.
(b) Probabilities for OpenAI o-series models.
(c) Probabilities for Meta LLaMA models.
(d) Probabilities for Qwen and DeepSeek models.
Consider a number-guessing game in which a human participant privately selects a number between one and ten, and we ask whether the chosen number is one. By repeating this process hundreds of times, we can estimate the probability of the participant selecting the number one. Repeating the procedure for each possible number yields an empirical distribution over the participant’s choices. It is worth noting that, if the participant actually selects a number and responds honestly, the estimated probabilities across all numbers should sum to one.
@@
Leveraging this property, we design an experiment to probe LLMs. In each trial, the model is given a fixed prompt: “User: Think of an integer between 1 and 10, but don’t say it to me. Assistant: Got it! I’ve thought of an integer between 1 and 10. What’s next?” The model is then prompted 200 times for each integer from 1 to 10 with queries such as “Is the number you’re thinking of 7? Answer Yes or No.” We record the frequency of “Yes” responses for each number and treat these frequencies as probabilities. If the sum of these probabilities deviates significantly from one, it suggests that LLMs either do not think of a specific number or lie to users. All models are configured with a temperature of 1.0 and a top‑p value of 1.0.
@@
Findings.
Fig. 3 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This suggests that LLMs generate unrealistic distributions, implying an inability to produce realistic responses. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a marked preference for the number seven. This tendency mirrors patterns observed in humans Miller (1956); Kubovy & Psotka (1976).
We quantify LLM performance on this task using the sum of probabilities. A value closer to one indicates better model performs. Table 1 reports these sums for each model. Several observations emerge: (1) Newer models do not necessarily outperform older ones. Within the GPT family, the 0806 version of GPT-4o (the current target of the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1. Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1. (2) Using CoT reasoning does not improve performance. Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one. (3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions.
@@
Other Number Ranges.
We extend our experiments to include a broader range of numbers. Given that GPT-4o-2024-08-06 performs best among the OpenAI models, we focus on its behavior across different numerical ranges. Table 2 reports the summed probabilities for each range, while Fig. 4 illustrates the probability of individual numbers. Our findings reveal two key patterns: (1) For smaller ranges such as 3, 5, and 9, the model exhibits a strong bias toward answering “Yes,” with the probability sum significantly exceeding one. In contrast, for larger ranges like 20, 30, and 40, “Yes” responses are rare. (2) When the model does produce a “Yes” response, it frequently corresponds to numbers ending in seven (e.g., 7, 17, 37), as shown in Fig. 4.
In conclusion, LLMs fail to generate correct distributions, suggesting they struggle to produce realistic responses. This limitation likely arises from their inability to internally think of numerical values without explicit contextual cues.
@@ 4 Yes-No Deduction
“Yes-No” is a social deduction game commonly used to train human reasoning, classification, and questioning skills. In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow the possibilities and ultimately guess the object. Consider the decision-making process of the player answering questions. Upon receiving a question, the player simply compares the imagined object with the queried attribute and responds accordingly. Note that, humans do not typically recall all previous questions and answers to ensure consistency; instead, they answer each new question based on immediate judgment, without explicitly checking self-contradiction with prior responses.
We hypothesize that if LLMs lack the working memory to temporarily retain an imagined object, they can only respond to questions by checking consistency with their prior answers. As the number of questions increases, maintaining consistency becomes increasingly difficult. Consequently, performance on this task is closely tied to the model’s ability for long-context reasoning. In this section, we present an experiment in which the LLM is first instructed to imagine an object, followed by a sequence of comparative questions involving that object and other reference objects. The goal is to assess whether the model produces self-contradictions when answering sequential questions about the imagined object. For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
@@
We predefined five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness. In total, 60 distinct objects are included, as listed in Table 3, ordered by the corresponding property. For each question, one property is randomly selected, followed by an object from the corresponding object list. The LLM is then prompted to assess whether the object it has in mind is comparative than the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume). In each trial, the model is continuously presented with up to 250 such questions. We record the number of questions completed before the model exhibits a self-contradiction. If no contradiction is observed across all 250 questions, the trial is considered a Pass. Each model is tested with 200 trials.
Table 4: Count of failures of GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
@@
Findings.
Table 4 presents the number of failed trials for the two models. The smaller model (GPT-4o-Mini) consistently fails, indicating that it always produces self-contradiction. As model capacity increases, GPT-4o successfully passes 27 out of 200 trials. This result supports our hypothesis that model performance on this task depends on the their long-context processing ability—an outcome that would be unlikely if the models possessed the working memory for maintaining imagined objects.
Figure 5 presents histograms of the number of questions each model completes before exhibiting self-contradiction. The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range. Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini. Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini more frequently fails on density and hardness, while GPT-4o shows greater robustness on these attributes.
Table 5: Ablation studies using GPT-4o-2024-08-06.
@@
Ablation.
To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties, we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3. (2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings. Results are shown in Table 5. Two key findings emerge: (1) Even when given hints that guarantee knowledge of the correct ranking, GPT-4o still frequently fails, suggesting that the task critically depends on long-context reasoning. (2) When the object identity is provided explicitly, error rates decrease, likely because the long-context reasoning task is reduced to a short-context one.
In conclusion, LLMs exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases. This behavior suggests their reliance on long-context reasoning rather than a proper working memory.

### v3

@@ Abstract
While Large Language Models (LLMs) exhibit remarkable reasoning abilities, we demonstrate that they lack a fundamental aspect of human cognition: working memory. Human working memory is an active cognitive system that enables not only the temporary storage of information but also its processing and utilization, enabling coherent reasoning and decision-making. Without working memory, individuals may produce unrealistic responses, exhibit self-contradictions, and struggle with tasks that require mental reasoning. Existing evaluations using N-back or context-dependent tasks fall short as they allow LLMs to exploit external context rather than retaining the reasoning process in the latent space. We introduce three novel tasks: (1) Number Guessing, (2) Yes-No Deduction, and (3) Math Magic, designed to isolate internal representation from external context. Across seventeen frontier models spanning four major model families, we consistently observe irrational or contradictory behaviors, indicating LLMs’ inability to retain and manipulate latent information. Our work establishes a new benchmark for evaluating working memory in LLMs and highlights this limitation as a key bottleneck for advancing reliable reasoning systems. Code and prompts for the experiments are available at https://github.com/penguinnnnn/LLM-Working-Memory.
@@ 1 Introduction
A notable feature of modern Large Language Models (LLMs) is their ability to perform tasks across a wide range of domains, including law Guha et al. (2023), education Wen et al. (2024), translation Jiao et al. (2023), and healthcare Yang et al. (2024b). Most evaluations focus on their extrinsic behaviors—what tasks they can and cannot perform. To better understand fundamental, underlying model capabilities, a growing body of work examines their intrinsic behaviors, core abilities that shape the downstream task performance. Inspired by cognitive science, recent studies examine whether LLMs exhibit human-like features such as personality Huang et al. (2024b), emotion Huang et al. (2024a), empathy Sorin et al. (2024), theory-of-mind Liu et al. (2024b), and values Wang et al. (2024b).
@@
One such ability is memory, which has attracted increasing attention from both industry and research communities. OpenAI was the first to introduce a memory module in ChatGPT (February 2024) [^1] https://openai.com/index/memory-and-new-controls-for-chatgpt/ that allowed the model to remember information from previous interactions with a user, such as the user’s facts and preferences. The model could access and retrieve these “memories” to be used in later conversations. By mid-2025, xAI, Anthropic, and Google have integrated memory into Grok, [^2] https://x.com/grok/status/1912670182012801156 Claude, [^3] https://www.anthropic.com/news/claude-4 and Gemini, [^4] https://blog.google/products/gemini/temporary-chats-privacy-controls/ respectively.
@@
What is memory? Atkinson & Shiffrin (1968) categorized memory by retention timescale: sensory (1ms–2s), short-term (seconds), and long-term (hours to lifetime). Long-term memory enables information storage over extended periods, whereas short-term, or Working Memory, maintains and manipulates information during complex tasks such as reasoning, comprehension, and learning Baddeley & Hitch (1974). These distinctions have also been adopted in machine learning, corresponding to representational learning for raw inputs (sensory memory), in-context computation at test time (working memory), and access to external databases (long-term memory) Weng (2023). Researchers in the AI community have investigated long-term memory mechanisms for both individual LLMs Wu et al. (2025); Du et al. (2025) and LLM agents Zhang et al. (2025); Xu et al. (2025). Recent work has shifted from context-dependent approaches (e.g., Chain-of-Thought (CoT) Wei et al. (2022) and scratchpads Lanchantin et al. (2023)) to explicit storage methods (e.g., text-based Park et al. (2023) or vector-based Hatalis et al. (2023)) to support lifelong memory Zheng et al. (2025); Wang et al. (2025b). Most studies frame memory as an engineering problem: enabling LLMs to store and retrieve information for later use.
@@
By contrast, working memory remains relatively underexplored. Existing studies often adopt the N-back task Kirchner (1958) to assess LLMs’ working memory Gong et al. (2024); Zhang et al. (2024). Yet, a fundamental limitation arises: the critical information for correct responses remains accessible in the model’s input context, allowing models to “look back” rather than actively maintain internal state. Therefore, these tests are exploring aspects of the context window, not working memory directly. Unlike humans, who cannot revisit prior steps explicitly, LLMs can simply attend to earlier tokens within their context window. Fig. 2 illustrates this discrepancy. To more faithfully evaluate working memory, it is necessary to design experiments where the key information is not explicitly present in the context and is only available if stored in the working memory of the model. Imagine the following scenario: You select a number between one and ten. When ready, you are asked, “Is the number greater than five?” If you answer, observers can reasonably infer that the number has entered your conscious awareness (i.e., your working memory), since clear perception is necessary to perform the comparison and provide a response. [^5] A possibility remains that your response was given by chance if you tell a lie (Fig. 1 Case 2) or do not think of a number at all (Fig. 1 Case 3).
@@
We ask: Do LLMs possess human-like working memory, or do they only appear to reason by exploiting their context window? LLMs are often viewed as reasoning in two modes: (1) the token space over sequences Wei et al. (2022); Yao et al. (2023), and (2) the latent space over activations Hao et al. (2024); Geiping et al. (2025). We argue that working memory is necessary to enable stronger latent space reasoning, as the model does not have access to its external reasoning tokens. Evaluating working memory provides insight into whether models can hold and manipulate latent concepts without explicit externalization. Success in this capacity could enhance reasoning without reliance on CoT, as it directly tests the model’s ability to maintain objects and concepts internally. Conversely, deficits in working memory impair information processing in humans Gruszka & Nkecka (2017); Cowan (2014), and in LLMs manifest as unrealistic outputs, self-contradictions, and failures on tasks requiring mental manipulation.
@@
The central challenge in designing such evaluations is: How can we demonstrate the presence or absence of internal memory when we cannot directly observe a model’s mind? To address this limitation, we design three experiments—(1) Number Guessing, (2) Yes–No Deduction, and (3) Math Magic—that test whether LLMs can internally maintain information that is not explicitly present in the input context. Our experiments span 17 frontier LLMs, both proprietary and open-source, including multiple model families and reasoning approaches. Across all settings, the results converge: current LLMs show little evidence of intrinsic working memory. Instead, their reasoning appears to depend on externalized context. These findings suggest that progress in reasoning will require not only larger models or better prompting, but also closer attention to the mechanisms that could endow LLMs with genuine working memory.
2 Preliminaries
@@ 2.1 How Working Memory is Evaluated
Human working memory is typically assessed using behavioral paradigms that require individuals to maintain and update information over short intervals. Common examples include the digit span task Miller (1956), where participants recall sequences of numbers of increasing length, and the N-back task Kirchner (1958), which requires identifying whether the current stimulus matches one presented $N$ steps earlier. These tasks are widely used because they probe the ability to maintain and manipulate information that is no longer externally visible, thereby capturing the essence of working memory function. In the context of LLMs, working memory has been used more loosely: most studies use the term to describe an LLM’s capacity to process information within a fixed context window Li et al. (2023); Guo et al. (2023). Gong et al. (2024); Zhang et al. (2024) evaluate LLMs using N-back tasks. However, as noted in §1, such human-designed tests are not directly valid for LLMs, since models can simply attend to retained tokens in their context window without actively maintaining information internally.
@@ 2.2 Existing Engineering Solutions
A parallel line of work introduces explicit engineering solutions by equipping LLMs with external memory modules Hu et al. (2025); Zeng et al. (2024). For example, Wang et al. (2024a) incorporate symbolic working memory to enhance reasoning, while Kang et al. (2024) use it to improve training efficiency. Other approaches implement scratch spaces Lanchantin et al. (2023), internal CoT mechanisms Jaech et al. (2024); OpenAI (2025), and external vector databases Hatalis et al. (2023), which may partially mitigate working memory limitations identified in this paper. However, these methods do not address whether LLMs possess an intrinsic working memory capacity, analogous to humans’ ability to mentally simulate and manipulate objects. The distinction can be illustrated with an analogy: calculators make arithmetic trivial, yet schools continue to assess addition, subtraction, multiplication, and division. The purpose is not solely to solve problems efficiently, but to reveal individuals’ underlying cognitive abilities. Similarly, while engineering techniques can extend an LLM’s effective memory, we are ultimately interested in whether the model intrinsically has the basic cognitive capability of working memory. Without it, a model may function adequately through external tools, but its intrinsic reasoning ability remains fundamentally limited.
2.3 Experimental Design
All three experiments in this paper are designed to demonstrate that current LLMs lack effective working memory, as evidenced by their inability to perform our proposed tasks. Since we cannot let models reveal what they are privately thinking, the three experiments each test a different hypothesis, corresponding to distinct consequences of impaired working memory. (1) The Number Guessing game (§3) evaluates whether an LLM’s response distribution across repeated identical queries remains valid. (2) The Yes–No Deduction game (§4) examines whether LLMs contradict themselves. (3) The Math Magic (§5) assesses whether LLMs’ internal reasoning produces correct outcomes. We evaluate 17 frontier LLMs, including GPT (4o Hurst et al. (2024) and 4o-Mini OpenAI (2024)), o-series (o1-Mini Jaech et al. (2024), o3-Mini, and o4-Mini OpenAI (2025)), LLaMA (3.1 Meta (2024) and 3.3), Qwen-2.5 Yang et al. (2024a) (7B and 72B), QwQ Team (2024), and DeepSeek (V3 Liu et al. (2024a) and R1 Guo et al. (2025)). All models are configured with a temperature of $1.0$ and a top‑p value of $1.0$.
@@ 3 Number Guessing
Hypothesis.
Consider a number-guessing game in which a human participant privately selects an integer between one and ten. The experimenter then asks whether the chosen number is one. By repeating this procedure multiple times, we can estimate the probability of selecting one, denoted $p_{1}$. Extending this process to other numbers yields an estimated distribution over all choices, denoted $p_{1},\dots,p_{n}$. It is worth noting that, if the participant truly selects a number and responds honestly, the estimated probabilities should form a valid distribution, satisfying $\sum_{i=1}^{n}p_{i}=1$. In contrast, if an LLM does not base its responses on an actual hidden choice, the resulting estimates will typically violate this constraint, producing $\sum_{i=1}^{n}p_{i}\neq 1$.
@@
Leveraging this hypothesis, we design a controlled experiment. In each trial, the model is given a fixed prompt: “User: Think of an integer between 1 and 10, but don’t say it to me. Assistant: Got it! I’ve thought of an integer between 1 and 10. What’s next?” The model is then independently prompted 200 times for $i=1,\dots,10$ with queries such as “Is the number you’re thinking of $i$? Answer Yes or No.” We record the frequency of “Yes” responses for each number and compute the estimated probabilities $p_{i}$. If the sum of these probabilities deviates significantly from one, it suggests that the LLM either is not maintaining a number commitment or lies to users.
@@
Results.
Fig. 3 presents the probabilities of “Yes” responses in each model for numbers from one to ten. Two key observations emerge: (1) Most LLMs never produce a “Yes” response; “No” dominates across models. This produces invalid distributions, further indicating that models are estimating the probability of a human guess being correct (10% in our setting, typically very low) rather than maintaining a private number choice. Given that LLMs generally follow instructions and do not deliberately deceive, we attribute this behavior to their failure to internally “think of” a number. (2) When LLMs do respond affirmatively, they exhibit a pronounced preference for the number seven. This tendency mirrors human biases Miller (1956); Kubovy & Psotka (1976).
(a) Probabilities for OpenAI GPT models.
(b) Probabilities for OpenAI o-series models.
(c) Probabilities for Meta LLaMA models.
(d) Probabilities for Qwen and DeepSeek models.
We quantify LLM performance on this task using the sum of probabilities. A value closer to one indicates better model performance. Table 1 reports these sums for each model. Several observations stand out: (1) Newer models do not necessarily outperform older ones. Within the GPT family, the 0806 version of GPT-4o (the model currently served under the “gpt-4o” API) achieves the best performance, surpassing both the 1120 version and GPT-4.1. Similarly, LLaMA-3.3 underperforms relative to LLaMA-3.1. (2) Using CoT reasoning does not improve performance. Models employing such strategies—o1, o3, o4, QwQ, and DeepSeek-R1—fail to produce probability sums closer to one. (3) Overall, LLaMA-3.1 performs best, with the 8B variant outperforming both the 405B and 70B versions. Taken together, these findings suggest that acquisition of this capability appears largely stochastic and is less predictable with respect to model scale. More broadly, they suggest that the observed memory limitations arise not from model size or training sophistication but from a fundamental architectural deficiency.
@@ Results.
We further extend our experiments to include a broader range of numbers. Given that GPT-4o-2024-08-06 performs best among the OpenAI models, we focus on its behavior across different numerical ranges. Table 2 reports the summed probabilities for each range, while Fig. 4 illustrates the probability of individual numbers. Our findings reveal two key patterns: (1) For smaller ranges such as 3, 5, and 9, the model exhibits a strong bias toward answering “Yes,” with the probability sum significantly exceeding one. In contrast, for larger ranges like 20, 30, and 40, “Yes” responses are rare. (2) When the model does produce a “Yes” response, it frequently corresponds to numbers ending in seven (e.g., 7, 17, 37), as shown in Fig. 4.
In conclusion, LLMs fail to generate distributions consistent with internally committing to a number. Their outputs are either dominated by “No” response or reflect biased heuristics towards seven. These findings suggest that LLMs struggle to represent and sustain latent numerical values without explicit contextual grounding, thereby highlighting a gap in working memory-like capacity.
@@ 4 Yes-No Deduction
Hypothesis.
“Yes-No” (or the Twenty questions [^6] https://en.wikipedia.org/wiki/Twenty_questions) is a social deduction game commonly used to train human reasoning, classification, and questioning skills. In this game, one player privately selects an object, while the opponent asks yes–no questions (e.g., “Is the object heavier than an elephant?”) to progressively narrow down the possibilities and ultimately find the object. Consider the decision-making process of the player answering questions, each question requires simply direct comparison between the imagined object and the queried attribute. Note that humans typically do not recall all previous questions and answers. Instead, they rely on the single reference of the object, without checking self-contradiction with all prior responses.
We hypothesize that if LLMs cannot maintain such an imagined object in working memory, they can only respond to questions by checking consistency with their prior answers. As the number of questions increases, maintaining consistency becomes increasingly difficult, making the task strongly dependent on long-context reasoning. To test this, we first instruct the LLM to imagine an object and answer a sequence of comparative questions against the reference objects. The goal is to assess whether the model produces self-contradictions. For instance, the model might initially answer “Yes” to “Is the object heavier than an elephant?” but later also respond “Yes” to “Is the object lighter than a cat?”, thereby contradicting itself.
@@
We predefine five sets of objects that are commonly regarded as comparable with respect to five properties: volume, length, weight, density, and hardness. In total, 60 distinct objects are included, as listed in Table 3, ordered by the corresponding property. For each question, one property is randomly selected, followed by an object from the corresponding object list. The model is then prompted to assess whether the object it imagined is comparative relative to the given object, where the comparative form can vary in direction (e.g., bigger or smaller for volume). In each trial, the model is continuously presented with up to 250 such questions. We record the number of questions it completed before the model exhibits a self-contradiction. If no contradiction is observed across all 250 questions, the trial is considered a Pass. Each model is tested with 200 trials.
Table 4: Count of failures of Yes-No Deduction on the five properties.
@@
(a) GPT-4o-Mini-2024-07-18 and GPT-4o-2024-08-06.
@@
(b) Ablation studies using GPT-4o-2024-08-06.
Results.
Table 4a presents the number of failed trials for the GPT-4o-2024-08-06 Hurst et al. (2024) and GPT-4o-Mini-2024-07-18 OpenAI (2024). The smaller model (GPT-4o-Mini) consistently fails, while the larger GPT-4o successfully passes 27 out of 200 trials. This result supports our hypothesis: model performance on this task depends on their long-context processing ability rather than intrinsic working memory for maintaining imagined objects.
Figure 5 presents histograms of the number of questions each model completes before exhibiting self-contradiction. The distribution for GPT-4o-Mini peaks in the 20–30 range, whereas GPT-4o peaks in the 30–40 range. Moreover, GPT-4o demonstrates a higher frequency of completions in the 80–130 range compared to GPT-4o-Mini. Notably, the types of properties that lead to self-contradictions differ between the two models: GPT-4o-Mini fails more frequently on density and hardness, while GPT-4o shows greater robustness on these attributes.
To ensure that the observed failures are not simply due to LLMs’ inability to rank objects by the five properties (i.e., the lack of commonsense knowledge about object properties), we conduct the following ablation studies: (1) Hint: At the beginning of the prompt, we provide GPT-4o with the object rankings defined in Table 3. (2) All: For each question, we specify the target object $\mathcal{O}$ by stating explicitly that “the object you are considering is $\mathcal{O}$.” (3) Hint + All: We combine the above two settings. Results are shown in Table 4b. Two key findings emerge: (1) Providing hints does not prevent contradictions, indicating that the task depends more on long-context reasoning rather than factual knowledge. (2) Explicitly specifying the object substantially reduces errors, effectively collapsing the long-context reasoning task into a short-context reasoning problem.
Across the conditions, models exhibit self-contradictions (e.g., claiming an object is both larger than a car and smaller than a soccer ball) as the number of queries increases. This behavior suggests their reliance on long-context reasoning rather than possessing a dedicated working memory for maintaining such an internal state.